Johan also decides that he wants to save for his retirement. He opens an investment account, earning 11% per annum compounded monthly.He decides to deposit R600 monthly for 40 years. What is the value of his investment at the end of the 40 year period


1
Expert's answer
2020-08-18T17:57:58-0400


this is an annuity


we solve by the formula

"FV=A\times\frac{(1+r)^n-1}{r}"


"r=\frac{0.11}{12}=0.0092"

"n=40\times12=480"

"FV=600\times\frac{(1+0.009)^{480}-1}{0.009}=5 160 076.32"